OpenStudy (anonymous):

The legs of a right triangle are 5m and 12m long. What is the length of the hypotenuse? Explain how you found your answer.

OpenStudy (anonymous):

According to Pythagorean Theorem, the hypo square = the sum of the legs' squares. So, hypo square = (5^2) + (12^2) , hypo square = 25 + 144 = 169. Therefore the hypo is the square root of 169, which is 13.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

U're very welcome. :)

OpenStudy (anonymous):

I apreccate it :)

OpenStudy (anonymous):

Oh great, how happy to hear that..

OpenStudy (anonymous):

Um I spelled that wrong but....yeah

OpenStudy (anonymous):

Its alrite, i understand. You're studying Pythagorean Theorem ?

OpenStudy (anonymous):

Yeah, Segment Exam. Very important to double check.